OpenStudy (diamondanon2):

Given: F(x) = 2x - 1; G(x) = 3x + 2; H(x) = x 2 Find F[G(x)] - F(x).

OpenStudy (diamondanon2):

Answer choices are 4x+4 4x+2 4x

OpenStudy (diamondanon2):

@jhonyy9

jhonyy9 (jhonyy9):

check it please there given again H(x) too but you dont wrote again used it i think you missed it from there yes ?

OpenStudy (diamondanon2):

F(x) = 2x - 1; G(x) = 3x + 2; H(x) = x^2

jhonyy9 (jhonyy9):

yes i understand it but on the find case there missed the H(x) where you need use it ?

OpenStudy (diamondanon2):

Oh I do not think you do. The find case is F[G(x)] - F(x)

OpenStudy (diamondanon2):

Is the question incorrect? Or can we ignore the H?

jhonyy9 (jhonyy9):

but than you need finding just this F(G(x)) -F(x) so why was given H(x) ?

OpenStudy (diamondanon2):

I do not know. Here is an image of the question.

jhonyy9 (jhonyy9):

ok. so than you are sure i accept it so

OpenStudy (diamondanon2):

I am sure.

jhonyy9 (jhonyy9):

ok i see it now but i think there is wrote wrong the last F(x) need being H(x) my opinion

OpenStudy (diamondanon2):

I see. What you are saying does make sense. So if we do it without H will be get an available answer?

jhonyy9 (jhonyy9):

common please try it so how wrote there and we will see what will result - ok. ?

OpenStudy (diamondanon2):

okay

jhonyy9 (jhonyy9):

so f(x) = 2x-1 g(x) = 3x+2 f(g(x)) = ?

jhonyy9 (jhonyy9):

using what i explained you before

OpenStudy (diamondanon2):

I got 4x+4

jhonyy9 (jhonyy9):

just substitute the g(x) in place of x inside equation of f(x) right ?

jhonyy9 (jhonyy9):

i think not is right please write me here step by step

OpenStudy (diamondanon2):

Oh no I mean for that i got 5x

OpenStudy (diamondanon2):

so 3x+2(x) = 5x

jhonyy9 (jhonyy9):

f(x) = 2x-1 g(x) = 3x+2 f(g(x)) = f(3x+2) = ?

jhonyy9 (jhonyy9):

in place of x substitute the 3x+2

jhonyy9 (jhonyy9):

f(x) = 2x-1 f(3x+2) = 2(3x+2) -1 = ?

OpenStudy (diamondanon2):

so it would be 6x+3

jhonyy9 (jhonyy9):

yes exactly - this is easy - yes ?

OpenStudy (diamondanon2):

yes is that the answer?

OpenStudy (diamondanon2):

or more steps?

jhonyy9 (jhonyy9):

yes

OpenStudy (diamondanon2):

okay :)

jhonyy9 (jhonyy9):

6x+3 so you can factorizing out the 3 and will get 3(2x+1)

jhonyy9 (jhonyy9):

this is the answer how is wrote there but i think there missed anything

jhonyy9 (jhonyy9):

this H(x) need used

OpenStudy (diamondanon2):

oh so thatt means the question is incorrect right?

jhonyy9 (jhonyy9):

exactly

OpenStudy (diamondanon2):

okay thank you a bunch!

jhonyy9 (jhonyy9):

do you have again ? and now do you understand it sure right ?

OpenStudy (diamondanon2):

Yes I am done. I understand. Thank you.

jhonyy9 (jhonyy9):

excouzme what mean this bunch ? bc i dont meet it till today

jhonyy9 (jhonyy9):

or this is just a slent used by students ?

jhonyy9 (jhonyy9):

sorry but i m from Europe and i learned alone English from books and internet and i dont know again this word of bunch ?

jhonyy9 (jhonyy9):

can you explain it me please

jhonyy9 (jhonyy9):

ty

jhonyy9 (jhonyy9):

bunch = much

jhonyy9 (jhonyy9):

ok. was my pleasure and than you wan,like anytime than i m here online just tag me there and i will try help you good luck bye bye

jhonyy9 (jhonyy9):

or do you wan we trying it with the last term to be H(x) ? probably in this way we will get an answer from given choices

jhonyy9 (jhonyy9):

Diamond are you here please ?